# Fact :Teaching MLLMs with Faithful, Concise and Transferable Rationales

### v1

@@ 4.3. Migrating to Large Models
We posit that CoT rationales generated by the student model could mutually benefit the teacher model as well. To test this theory, we devised an experimental framework that applied two specific types of contextual prompts to the mPlug-Owl 13B (Ye et al., 2024) model in the GQA and OK-VQA context: 1) a direct question followed by its answer, and 2) a question accompanied by CoTs produced by task-specific models, plus the answer, with a consistent presentation of three prompts. The outcomes observed on both tasks on test sets are compiled in Table 2. These findings corroborate our assertion that CoTs do indeed bolster the inferential precision of teacher models. Interestingly, variations in performance enhancements attributed to CoTs sourced from MiniGPT4 and OpenFlamingo were minimal when applied to the larger model. This phenomenon likely stems from the extensive comprehension capabilities inherent to larger models, enabling them to effectively leverage the supplied CoTs for detailed and accurate deductions. This not only validates the utility of CoTs across model scales but also highlights the adaptability of larger models to assimilate and refine input from smaller counterparts, further emphasizing the symbiotic potential between models of differing capacities.
Table 4. Ablation results (%) of individual components.
GQA OK-VQA
0 Backbone 30.1 28.2
1    + faithfulness 30.8 29.7
2    + conciseness 31.9 31.5
3    + transferability 32.4 31.8
@@ 4.4. Ablation Experiment
Qualitative Analysis. We begin by presenting several qualitative examples in Figure  4 to illustrate our process from visual programming to symbolic trace, and subsequently to multimodal CoT rationales. Throughout progression, we continually refine the capabilities of large models into more comprehensible language, ultimately distilling this knowledge into MLLMs. This approach demonstrates the efficacy of our method in harnessing the sophisticated reasoning abilities of teacher models and making them accessible to MLLMs through a distilled, understandable format.
Analysis on Fact. For the three properties of CoT proposed in our paper, we conducted sequential ablation experiments using OpenFlamingo (3B) as the backbone, with results on the GQA documented in Table  4: 1) Backbone: Merely use a language model to convert the entire execution trace into CoTs for distillation. 2) Backbone + faithfulness: Here, we filter the program and retain pieces of code that produce the correct answers, thereby ensuring faithfulness. 3) Backbone + faithfulness + conciseness: At this stage, operations such as pruning, merging, and bridging are performed during execution to enhance conciseness, eliminating redundancy. 4) Backbone + faithfulness + conciseness + transferability: Finally, we focus on transferability, selecting rationales that are most suitable for distillation. The results demonstrate that each of the three properties significantly contributes to performance enhancement. Faithful CoTs enable the model to infer correct outcomes, conciseness helps the model focus more on reasoning logic, and transferability effectively filters data suited for distillation, achieving a more suitable knowledge representation for smaller models.
@@
Human Evaluation We manually chose 100 responses from the GQA dataset and manually undertook a detailed error analysis. The errors identified were classified into four main categories: Logical, Factual, Format, and Localization errors. We will detail the types of errors in Appendix C. For MiniGPT4, we utilized the prompt, ”Answer the following VQA questions step by step.” Subsequently, we juxtaposed these outcomes with those derived from using Fact, illustrating the comparative analysis in a graphically represented figure  6. A noteworthy observation is a pronounced decrease in both Logical errors and Localization errors. This is consistent with our expectation that Fact can improve the reasoning and spatial capabilities of MLLMs. However, we conjecture that constraints inherent to the model parameters contributed to the persistence of Format errors across both methodologies. This suggests that while our approach significantly mitigates certain types of errors, the challenge of completely eliminating format-related inaccuracies remains, indicating a potential area for further refinement in model training and prompt engineering strategies.
@@ 5. Conclusion
This study presents Fact for generating, refining, and distilling CoT rationales to enhance multimodal models’ reasoning capabilities. Through targeted CoT editing operations such as pruning, merging, and bridging, we efficiently remove irrelevant information and improve the coherence of CoTs. By teaching these high-quality rationales to MLLMs, we significantly boost model performance across various tasks. Our experimental results underscore the importance of precise CoT rationale refinement, demonstrating marked performance enhancements in both teacher and student models.

### v2

@@ 4.2. Quantitative Results
GQA OK-VQA
mPlug-Owl (13B) 56.5 57.6
+ CoT (GPT-4v) 62.5 62.1
+ CoT (OpenFlamingo-Fact) 62.7 62.6
MiniGPT4 (7B) 34.5 35.1
+ CoT (GPT-4v) 38.4 40.8
+ CoT (OpenFlamingo-Fact) 38.9 41.3
@@ 4.3. Migrating to Large Models
We posit that CoT rationales generated by the student model could mutually benefit the teacher model as well. To test this theory, we devised an experimental framework that applied two specific types of contextual prompts to the mPlug-Owl 13B (Ye et al., 2024) model in the GQA and OK-VQA context: 1) a direct question followed by its answer, and 2) a question accompanied by CoTs produced by task-specific models (GPT-4v and OpenFlamingo-Fact), plus the answer, with a consistent presentation of three prompts. The outcomes observed on both tasks on test sets are compiled in Table 3. These findings corroborate our assertion that CoTs do indeed bolster the inferential precision of teacher models and OpenFlamingo 3B using Fact generates better CoT than that generated by GPT-4v, demonstrating the Fact’s superiority in generating high-quality CoT. This not only validates the utility of CoTs across model scales but also highlights the adaptability of larger models to assimilate and refine input from smaller counterparts, further emphasizing the symbiotic potential between models of differing capacities.
Table 4. Ablation results of individual components.
VQAv2 GQA OK-VQA TallyQA MME MM-Bench SEED
Simple Complex
Backbone 44.6 30.1 28.2 64.4 59.3 668.2 13.6 33.1
+faithfulness 47.3 30.8 29.7 66.9 63.6 884.5 16.0 37.4
+conciseness 48.7 31.9 31.5 68.7 64.9 903.4 17.3 38.6
+transferability 49.2 32.4 31.8 70.1 65.7 912.2 18.3 39.2
Instruct generalist 45.9 30.9 30.3 65.9 61.8 847.3 14.8 34.9
VPD 46.8 30.8 30.2 66.6 62.5 870.7 15.2 35.8
@@ 4.4. Ablation Experiment
Qualitative Analysis. We begin by presenting a qualitative example in Figure  4 to illustrate our process from visual programming to symbolic trace, and subsequently to multimodal CoT rationales. Throughout progression, we continually refine the capabilities of large models into more comprehensible language, ultimately distilling this knowledge into MLLMs. This approach demonstrates the efficacy of our method in harnessing the sophisticated reasoning abilities of teacher models and making them accessible to MLLMs through a distilled, understandable format.
Analysis on Fact. For the three properties of CoT proposed in our paper, we conducted sequential ablation experiments using OpenFlamingo (3B) as the backbone, with results on the GQA documented in Table  4: 1) Backbone: Merely use a language model to convert the entire execution trace into CoTs for distillation. 2) Backbone + faithfulness: Here, we filter the program and retain pieces of code that produce the correct answers, thereby ensuring faithfulness. 3) Backbone + faithfulness + conciseness: At this stage, operations such as pruning, merging, and bridging are performed during execution to enhance conciseness, eliminating redundancy. 4) Backbone + faithfulness + conciseness + transferability: Finally, we focus on transferability, selecting rationales that are most suitable for distillation. The results demonstrate that 1) incorporating faithfulness enhances model’s performance on various tasks, surpassing that of instruction tuning. The minor discrepancies in some datasets can be attributed to the redundancy and transferability challenges that hinder the model’s learning process. These issues are completely resolved after integrating conciseness and transferability. 2) the Fact paradigm markedly outperforms both the instruct tuning and VPD highlights the Fact’s effectiveness in high-quality CoT generation. 3) Fact employs a more fine-grained approach to analyze and edit code, addressing the issues of redundancy and logical gaps that arise from VPD’s direct conversion of code into CoT.
@@
Human Evaluation. We manually chose 100 responses from the GQA dataset and manually undertook a detailed error analysis. The errors identified were classified into four main categories: Logical, Factual, Format, and Localization errors. We will detail the types of errors in Appendix C. For MiniGPT4, we utilized the prompt, ”Answer the following VQA questions step by step.” Subsequently, we juxtaposed these outcomes with those derived from using Fact, illustrating the comparative analysis in a graphically represented figure  5. A noteworthy observation is a pronounced decrease in both Logical errors and Localization errors. This is consistent with our expectation that Fact can improve the reasoning and spatial capabilities of MLLMs. However, we conjecture that constraints inherent to the model parameters contributed to the persistence of Format errors across both methodologies. This suggests that while our approach significantly mitigates certain types of errors, the challenge of completely eliminating format-related inaccuracies remains, indicating a potential area for further refinement in model training and prompt engineering strategies.
@@ 5. Conclusion
This study presents Fact for generating, refining, and distilling CoT rationales to enhance multimodal models’ reasoning capabilities. Through targeted CoT editing operations, we remove irrelevant information and improve the coherence of CoTs. By teaching these high-quality rationales to MLLMs, we boost model performance across various tasks. Our experimental results underscore the importance of precise CoT rationale refinement, demonstrating marked performance enhancements in both teacher and student models.
6. Acknowledgment
This work was supported by the National Key Research and Development Project of China (2022ZD0160101), Key Research and Development Projects in Zhejiang Province (No. 2024C01106), the NSFC (No. 62272411), and Research funding from FinVolution Group.